OpenStudy (anonymous):

Us Intermediate Value theorem to show that F(x)=sin(x)=1 has a zero

OpenStudy (zzr0ck3r):

sin(x) = 1 is not a function

OpenStudy (zzr0ck3r):

did you mean sin(x)+1?

OpenStudy (zzr0ck3r):

that would not make sense either actually...please edit the question.

OpenStudy (freckles):

f(x)=sin(x)+1 does have a zero when sin(x)=-1 so why would f(x)=sin(x)+1 not make sense?

OpenStudy (freckles):

though he could have meant also f(x)=sin(x)-1 and that would have zeros when sin(x)=1

OpenStudy (freckles):

I wonder if also there was a meant to be some interval given

jimthompson5910 (jim_thompson5910):

If you meant to say f(x) = sin(x)+1, you cannot use the intermediate value theorem. Why not? Because there is no sign change (sin(x)+1 is always nonnegative) needed to show there is a root.

OpenStudy (freckles):

oh true never negative

OpenStudy (zzr0ck3r):

right that is what I was referring to when I said it would not make sense. Continuous on [a,b] and we get an element in (a,b), so 1 or -1 would not make sense.... no idea